# Joint Reconstruction of Spatially-Coherent and Realistic Clothed Humans and Objects from a Single Image

### v1

@@ 4 Experiments
Datasets: Due to the lack of 3D human-object datasets, we introduce the dataset synHOR for training our model. synHOR is a synthetic dataset created using $526$ 3D human scans from Thuman 2.0 [44] and $27$ 3D object scans including all 20 objects from BEHAVE [2] and selected 7 from HODome [46]. We picked random pairs of human and object meshes, where the object was initialized with a random pose and optimized to be in contact with the human. We then simulated 6 random translations of the human-object pair within the FOV of a perspective camera placed at the origin, and rendered 180 views for each translation. Finally, we discard the images where the object is not in view. We train SCHOR with 500 subjects of synHOR. For quantitative evaluation, we use 99 images from synHOR. We test the generalization power of SCHOR by inferring the human-object shapes from real images taken from BEHAVE [2]. BEHAVE is a real-world dataset that captures interactions between 8 humans and 20 objects. Due to its real-world setting, high-quality ground truth meshes are unavailable; therefore, we perform only a qualitative evaluation.
@@ 4.1 Ablation Study
Effectiveness of architecture configuration: We demonstrate the significant improvement achieved by using the architecture configuration explained in Sec. 3.2. 6 alternate configurations of the architecture of the attention-based neural implicit model of SCHOR are trained and tested:
  • Single MLP: To prove the effectiveness of estimating an implicit representation for each human and object, this configuration uses a single MLP to estimate a single occupancy $\hat{s}$ for both human and object.
@@
Both quantitative (Tab. 1) and qualitative (Fig. 4) results prove that the SCHOR’s architecture outperforms the other configurations. Specifically, replacing two MLPs with one (Single MLP) reduces reconstruction accuracy and detail, as seen in the hand of the BEHAVE model. If two MLPs process the same feature obtained by merging all the features with a single transformer (Single Trans), human-specific information is lost, resulting in an incomplete human reconstruction. Using both a single MLP and a single transformer degrades quality, producing smoother surfaces than SCHOR. Omitting the transformer encoder altogether (No Trans) makes the network less robust and introduces noise in the reconstruction. Finally, concatenating the input image $I_{f}$ directly with $I_{o}$ and $I_{h}$, significantly reduces quality compared to SCHOR, both when a transformer is used to merge human-object features (Concat Trans) and when it is not (Concat No Trans), proving that extracting features directly from $I_{f}$ improves performance.
@@
Effect of different inputs: SCHOR uses multiple input data to reconstruct the human-object shape, including the RGB image $I_{f}$, full-body human image $I_{h}$, 2D surface normal $S_{N}$, segmented object $I_{o}$ and human-object pose. In this ablation study, we show the advantages of combining all these inputs. We adapt SCHOR architecture to process different combinations of the input data. In each configuration, two MLPs are still used to estimate $\hat{s}_{h}$ and $\hat{s}_{o}$.
  • No $\mathbf{S_{N}}$: Same configuration as SCHOR but without 2D normal map $S_{N}$ concatenated to $I_{h}$.
  • No $\bm{\sigma_{\{h,o\}}}$: Same configuration as SCHOR but without processing the human-object pose features $\sigma_{\{h,o\}}$.
  • Only RGB: Same configuration as SCHOR excluding 2D normal map $S_{N}$ and human-object pose features $\sigma_{\{h,o\}}$.
@@
synHOR
PIFuHD 33.75 67.59 .4555 .6595 18.07
ECON 36.48 79.39 .3952 .6547 14.57
SiTH 31.88 68.88 .3371 .6521 14.33
$\mathrm{PIFuHD}_{\mathrm{ho}}$ 43.50 91.97 .3917 .5661 11.00
$\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$ 31.66 50.78 .4228 .7090 20.59
$2\mathrm{PIFuHD}_{\mathrm{ho}}$ 48.29 80.09 .3482 .5347 10.79
$2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$ 28.82 36.83 .6966 .7196 25.62
$2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$ 25.83 33.20 .8176 .7204 29.17
$2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$ 23.87 26.53 .8172 .7601 32.72
SCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 3: Quantitative comparisons on synHOR dataset between SCHOR and related works that aim to reconstruct high-quality 3D shape in the bottom part. Results from comparisons with baselines created from PIFuHD [29] for fair comparisons are also presented.
@@
The benefit of incorporating all input data types is demonstrated by the superior quantitative results in Tab. 2. This is further supported by the visual results in Fig. 5, where more realistic and accurate human-object shapes are obtained with SCHOR. When the input RGB image is omitted (No $\mathbf{I_{f}}$), the model lacks global scene understanding, resulting in severe artifacts. Likewise, if human and object images are not used (Only $\mathbf{I_{f}}$), the network lacks local information about them. Only RGB shows the importance of using both normal maps and pose features to reproduce realistic details and ensure 3D spatial coherency in the reconstruction. Additional ablations are in the supplementary.
Figure 7: Qualitative evaluations against methods which aim to model the human-object interaction with examples from BEHAVE dataset [2]. Note that HDM generates point clouds rather than meshes. Front and side views are shown.
4.2 Comparisons
Our goal is to demonstrate that SCHOR can jointly reconstruct 3D human-object shapes from single-view images, producing realistic, detailed, and spatially coherent outputs. To the best of our knowledge, SCHOR is the first framework to address this task since related works either focus solely on realistic human reconstruction or model human-object interactions using template-based shapes, limiting detail and reducing the realism of reconstructions. We compare SCHOR with approaches of both categories.
Realistic human reconstruction methods: We evaluate SCHOR against methods designed for high-quality human shape reconstruction from single images using neural implicit models (PIFuHD [29], ECON [41], SiTH [8]). To demonstrate that these works cannot jointly reconstruct human and object shapes, we use the RGB image $I_{f}$ as input. For a fair comparison, we retrain PIFuHD on synHOR ($\mathrm{PIFuHD}_{\mathrm{ho}}$) and introduce several baseline models based on the PIFuHD architecture to showcase SCHOR’s superiority in joint human-object reconstruction. We first use two MLPs instead of one to estimate an implicit representation of both human and object ($2\mathrm{PIFuHD}_{\mathrm{ho}}$). Since PIFuHD does not rely on 3D spatial information, we repeat the previous experiments by concatenating the human-object pose feature $\sigma$ to the extracted features (indicated with the prefix $\sigma$). Then, features are extracted from only the segmented human and object images ($2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$) and finally, the images ($I_{f},I_{h},I_{o}$) are concatenated together as input before feature extraction ($2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$).
As shown in Tab. 3, SCHOR achieves the best quantitative results and, as seen in Fig. 6, is the only approach able to reconstruct realistic, spatially coherent human-object shapes without noise. Human-focused approaches (ECON, SiTH, PIFuHD) struggle with object reconstruction, impacting human shape quality as well. Retraining these models with the synHOR dataset ($\mathrm{PIFuHD}_{\mathrm{ho}}$) and even adding two MLPs and pose parameters ($2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$) is insufficient to achieve the realism of SCHOR, which outperforms all approaches significantly.
Parametric human-object reconstruction methods: We also compare SCHOR against state-of-the-art parametric-model based methods for human-object reconstruction from a single image, mainly CHORE [36], CONTHO [20] and HDM [38]. As shown in  Fig. 7, SCHOR is the only method that reconstructs realistic clothed human-object shapes. Related works use a parametric-based representation of humans, which lacks detail and significantly reduces realism. These parametric approaches are designed to model human-object interactions explicitly. Our results show that despite not being trained on an explicit human-object interaction dataset, SCHOR implicitly reconstructs human-object interactions due to its spatial awareness. Note that we do not quantitatively evaluate parametric methods as the realism of their shapes is much lower than non-parametric shapes.
See supplementary for additional comparison results.
Limitations and Future work: Realistic clothed human-object reconstruction is an extremely challenging problem. In this work, we do not explicitly model human-object interactions, which can be less accurate than related works that focus on this. We specifically address the case where the human is occluded by an object, and not vice-versa. Finally, SCHOR leverages object poses estimated from known object templates, limiting its generalization to unseen objects. In future work, we will address these limitations by proposing a method that can handle occluded objects and explicitly model human-object interactions. We will also aim to retrieve textures of the clothed human-object reconstruction.
@@ 5 Conclusion
We presented SCHOR, the first novel method that can jointly reconstruct realistic clothed humans and objects from a human-object scene image. SCHOR combines the power of attention-based neural implicit learning with a generative diffusion model and human-object pose conditioning to produce detailed, spatially coherent human-object 3D shapes. Extensive ablations on various network architecture components of SCHOR demonstrate the effectiveness of the proposed approach. Our experiments show that our method generalizes well to a real world dataset on which it was not trained and outperforms the state-of-the-art methods in reconstructing realistic human-object scenes.
@@ 6 Overview
  • Additional details about the implementation of SCHOR (Sec. 7);
@@
  • Additional results of SCHOR and related works on both synHOR and BEHAVE [2] datasets (Sec. 10);
@@ 7.1 Diffusion model
Inspired by SitH [8], our image-conditioned latent diffusion model [26] used to generate missing body regions as described in Sec. 3.1, is trained by fine-tuning the diffusion U-Net’s [26] weights. These weights are initialized using the Zero-1-to-3 [18] model, combined with a trainable ControlNet [48] model, following the default network setups with input channel adjustment. The ControlNet input is a 1 channel mask and the diffusion U-Net input is a 3-channel RGB image of size 512 x 512. We train the models with a batch size of 6 images on a single A100 NVIDIA GPU, with a learning rate of $64\times 10^{-}$ 6 and adopting a constant warmup scheduling. The ControlNet model’s conditioning scale is fixed at 1.0. We use classifier-free guidance in our training, involving a dropout rate of 0.05 for the image-conditioning. During inference, a classifier-free guidance scale of 2.5 is applied to generate the output images.
@@ 7.2 Attention-based neural implicit model
where $QW^{q}_{i}$, $KW^{k}_{i}$,$VW^{v}_{i}$ are the parameters of $Q$, $K$ and $V$, and $W^{o}$o the parameters of the final projection. The final feature is computed as the mean of the processed features. with each merged feature $\varphi_{\{h,o\}}$ containing local information from the human or object image along with global scene information from the input image. Finally, to estimate the neural implicit representations, each decoder $f_{h}$, $f_{o}$ is an MLP with the number of neurons (259, 1024, 512, 256, 128, 1) and skip connections at 2nd, 3rd and 4th layers. We train the network end-to-end using Adam optimizer with a learning rate of $41e-$ 4 and batch size 4.
@@ 9.1 Effect of different inputs
Fig. 5 of the main paper shows examples of human-object shapes reconstructed using different combinations of input. Additional results from the configurations analyzed in the second ablation study of Sec. 4.1 and not included in Fig. 5 are shown in Fig. 8. The highest quality human-object shapes are obtained using SCHOR, confirming what was proved in Sec. 4.1. Smoother shapes are reconstructed when normal maps are excluded (No $S_{N}$) while omitting pose features (No $\sigma_{{h,o}}$) introduces depth ambiguity. Excluding the full-body human input (No $I_{h}$) or the object input (No $I_{o}$) prevents the network from reconstructing the human or object, respectively.
@@ 9.2 Training configurations
SCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 5: Quantitative results obtained with different training configurations of SCHOR.
Figure 9: Examples of results obtained with different training configurations of SCHOR. Front and side views are shown from synHOR in the top row, from BEHAVE [2] in the bottom.
The proposed attention-based neural implicit model is trained end-to-end to enable better joint reasoning about the human-object scene. This section presents results obtained by modifying SCHOR’s training configuration:
@@
As demonstrated by the quantitative evaluation in Tab. 5 and the qualitative results (Fig. 9), training the attention-based neural implicit model end-to-end significantly improves performance. Both $\mathbf{Sep}$ and $\mathbf{Sep_{no}}$ training configurations result in less accurate human-object reconstructions.
Figure 10: Effect of not applying the inpainting module of SCHOR before estimating the implicit representation. The same examples used in the paper are illustrated with (w.) and without (w.o) the object.

### v2

@@ 4 Experiments
Datasets: Due to the lack of 3D human-object datasets with high-quality meshes that represent clothing details, we introduce the dataset synHOR to train our model. synHOR is a synthetic dataset created using $526$ 3D human scans from Thuman 2.0 [42] and $27$ 3D object scans including all 20 objects from BEHAVE [2] and selected 7 from HODome [44]. We picked random pairs of human and object meshes, where the object was initialized with a random pose and optimized to be in contact with the human. We then simulated 6 random translations of the human-object pair within the FOV of a perspective camera placed at the origin, and rendered 180 views for each translation. Finally, we discard the images where the object is not in view. We train ReCHOR with 500 subjects of synHOR. For quantitative evaluation, we use 99 images from synHOR. We test the generalization power of ReCHOR by inferring the human-object shapes from real images taken from BEHAVE [2]. BEHAVE is a real-world dataset that captures interactions between 8 humans and 20 objects. Due to its real-world setting, high-quality ground truth meshes are unavailable; therefore, we perform only a qualitative evaluation.
@@
4.1 Comparisons
Our goal is to demonstrate that ReCHOR can jointly reconstruct realistic clothed human and object shapes from single images. To the best of our knowledge, ReCHOR is the first framework that represents realistic human details in the joint reconstruction using an unconstrained topology for the human-object shape since related methods rely on template-based or coarse representations, reducing reconstruction realism. We also compare ReCHOR with approaches that focus solely on high-quality human reconstruction to highlight their inability to reconstruct objects.
Figure 4: Qualitative evaluations against methods which aim to reconstruct human-object jointly with examples from BEHAVE dataset [2]. Note that HDM generates point clouds rather than meshes. Front and side views are shown.
Parametric human-object reconstruction methods: We compare ReCHOR against state-of-the-art human-object reconstruction methods from a single image that use a template-based or coarse representation of the 3D shapes, namely CHORE [34], CONTHO [18] and HDM [36]. As shown in  Fig. 4, ReCHOR is the only method that reconstructs realistic clothed human shapes along with the objects. Related works use a parametric-based representation of humans, which lacks detail and significantly reduces realism. These approaches are designed solely to predict the human-object spatial configuration. ReCHOR leverages them as priors and significantly improves the realism of the reconstruction while preserving the 3D spatial configuration. Note that we do not quantitatively evaluate parametric methods as the quality of their human shapes is much lower than non-parametric shapes.
synHOR
PIFuHD 33.75 67.59 .4555 .6595 18.07
ECON 36.48 79.39 .3952 .6547 14.57
SiTH 31.88 68.88 .3371 .6521 14.33
$\mathrm{PIFuHD}_{\mathrm{ho}}$ 43.50 91.97 .3917 .5661 11.00
$\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$ 31.66 50.78 .4228 .7090 20.59
$2\mathrm{PIFuHD}_{\mathrm{ho}}$ 48.29 80.09 .3482 .5347 10.79
$2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$ 28.82 36.83 .6966 .7196 25.62
$2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$ 25.83 33.20 .8176 .7204 29.17
$2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$ 23.87 26.53 .8172 .7601 32.72
ReCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 1: Quantitative comparisons on synHOR dataset between ReCHOR and related works that reconstruct high-quality 3D shape in the bottom part. Results from comparisons with baselines created from PIFuHD [27] are also presented.
Realistic human reconstruction methods: We evaluate ReCHOR against methods designed for high-quality human shape reconstruction from single images using neural implicit models (PIFuHD [27], ECON [39], SiTH [8]). To demonstrate that these works cannot jointly reconstruct human and object shapes, we use the RGB image $I_{f}$ as input. We also retrain PIFuHD on synHOR ($\mathrm{PIFuHD}_{\mathrm{ho}}$) and introduce several baseline models based on the PIFuHD architecture to showcase ReCHOR’s superiority in joint human-object reconstruction. We first use two MLPs instead of one to estimate an implicit representation of both human and object ($2\mathrm{PIFuHD}_{\mathrm{ho}}$). Since PIFuHD does not rely on 3D spatial information, we repeat the previous experiments by concatenating the human-object pose feature $\sigma$ to the extracted features (indicated with the prefix $\sigma$). Features are then extracted from only segmented human and object images ($2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$) and finally, the images ($I_{f},I_{h},I_{o}$) are concatenated together as input before feature extraction ($2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$). We do not repeat these experiments with ECON and SiTH because adapting their methods to our dataset would require substantial modifications. As shown in Tab. 1, ReCHOR achieves the best quantitative results and, as seen in Fig. 5, is the only approach that can jointly reconstruct realistic clothed human and object shapes without noise. Human-focused approaches (ECON, SiTH, PIFuHD) struggle with object reconstruction, impacting human shape quality as well. Retraining PIFuHD with synHOR dataset ($\mathrm{PIFuHD}_{\mathrm{ho}}$) and even adding two MLPs and pose parameters ($2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$) is insufficient to achieve the realism of ReCHOR, which outperforms all methods significantly. See supplementary for more comparisons.
@@ 4.2 Ablation Study
Effectiveness of architecture configuration: We demonstrate the significant improvement achieved by using the architecture configuration explained in Sec. 3.2. 6 alternate configurations of the architecture of the attention-based neural implicit model of ReCHOR are trained and tested:
  • Single MLP: To prove the effectiveness of estimating an implicit representation for each human and object, a single MLP is used on concatenated features $\varphi_{\{h,o\}}$ to estimate a single occupancy $\hat{s}$ for both human and object.
@@
Both quantitative (Tab. 2) and qualitative (Fig. 6) results prove that the ReCHOR’s architecture outperforms the other configurations. Specifically, replacing two MLPs with one (Single MLP) reduces reconstruction accuracy and detail, as seen in the hand of the BEHAVE model. If two MLPs process the same feature obtained by merging all the features with a single transformer (Single Trans), human-specific information is lost, resulting in an incomplete human reconstruction. Using both a single MLP and a single transformer degrades quality, producing smoother surfaces than ReCHOR. Omitting the transformer encoder altogether (No Trans) makes the network less robust and introduces noise in the reconstruction, showing the importance of global-local contextualization. Finally, concatenating the input image $I_{f}$ directly with $I_{o}$ and $I_{h}$, significantly reduces quality compared to ReCHOR, both when a transformer is used to merge human-object features (Concat Trans) and when it is not (Concat No Trans), proving that extracting features directly from $I_{f}$ improves performance.
@@
Effect of different inputs: ReCHOR uses multiple inputs to reconstruct the 3D human-object, including the RGB image $I_{f}$, full-body human image $I_{h}$, 2D surface normal $S_{N}$, segmented object $I_{o}$ and human-object poses $\sigma_{\{h,o\}}$. In this ablation study, we show the advantages of combining all these inputs. We adapt ReCHOR architecture to process different combinations of input data. In each configuration, two MLPs are still used to estimate $\hat{s}_{h}$ and $\hat{s}_{o}$.
  • No $\mathbf{S_{N}}$: Same as ReCHOR but without 2D normal map $S_{N}$ concatenated to $I_{h}$.
  • No $\bm{\sigma_{\{h,o\}}}$: Same as ReCHOR but without processing the human-object pose features $\sigma_{\{h,o\}}$.
  • Only RGB: Same as ReCHOR without 2D normal map $S_{N}$ and human-object pose features $\sigma_{\{h,o\}}$.
@@
The benefit of incorporating all input data types is demonstrated by the superior quantitative results in Tab. 3. This is further supported by the visual results in Fig. 7, where more realistic and accurate human-object shapes are obtained with ReCHOR. When the input RGB image is omitted (No $\mathbf{I_{f}}$), the model lacks global scene understanding, resulting in severe artifacts. Likewise, if human and object images are not used (Only $\mathbf{I_{f}}$), the network lacks local information about them. Only RGB shows the importance of using both normal maps and pose features to reproduce realistic details and avoid depth ambiguity in the reconstruction. Additional ablations are in the supplementary.
Limitations and Future work: Modeling realistic clothed human and object interactions is an extremely challenging problem. This work represents the first step toward achieving this goal. The accuracy of the interaction between humans and objects will be further improved in future works. We also aim to improve the quality of the object shape by introducing higher quality ground-truth object shapes in the dataset to learn high-frequency details for the objects as well. While our approach specifically addresses cases where a human is occluded by an object, we will extend it to handle object occlusions as well. We also aim to retrieve textures for clothed human-object reconstructions.
@@ 5 Conclusion
We presented ReCHOR, the first novel method that can jointly reconstruct realistic clothed humans and objects from a human-object scene image. ReCHOR combines the power of attention-based neural implicit learning with a generative diffusion model and human-object pose conditioning to produce realistic clothed human and object 3D shapes. Extensive ablations on various network architecture components of ReCHOR demonstrate the effectiveness of the proposed approach. Our experiments show that our method generalizes well to a real world dataset on which it was not trained and outperforms the state-of-the-art methods in reconstructing realistic human-object scenes.
@@ 6 Overview
  • Additional details about the implementation of ReCHOR (Sec. 7);
@@
  • Additional results of ReCHOR and related works on both synHOR and BEHAVE [2] datasets (Sec. 10);
@@ 7.1 Diffusion model
Inspired by SitH [8], our image-conditioned latent diffusion model [24] used to generate missing body regions as described in Sec. 3.1, is trained by fine-tuning the diffusion U-Net’s [24] weights. These weights are initialized using the Zero-1-to-3 [lugaresi2019mediapipe] model, combined with a trainable ControlNet [46] model, following the default network setups with input channel adjustment. The ControlNet input is a 1 channel mask and the diffusion U-Net input is a 3-channel RGB image of size 512 x 512. We train the models with a batch size of 6 images on a single A100 NVIDIA GPU, with a learning rate of $64\times 10^{-}$ 6 and adopting a constant warmup scheduling. The ControlNet model’s conditioning scale is fixed at 1.0. We use classifier-free guidance in our training, involving a dropout rate of 0.05 for the image-conditioning. During inference, a classifier-free guidance scale of 2.5 is applied to generate the output images.
@@ 7.2 Attention-based neural implicit model
where $QW^{q}_{i}$, $KW^{k}_{i}$,$VW^{v}_{i}$ are the parameters of $Q$, $K$ and $V$, and $W^{o}$o the parameters of the final projection. The final feature is computed as the mean of the processed features with each merged feature $\varphi_{\{h,o\}}$ containing local information from the human or object image along with global scene information from the input image. Finally, to estimate the neural implicit representations, each decoder $f_{h}$, $f_{o}$ is an MLP with the number of neurons (259, 1024, 512, 256, 128, 1) and skip connections at 2nd, 3rd and 4th layers. We train the network end-to-end using Adam optimizer with a learning rate of $41e-$ 4 and batch size 4.
@@ 9.1 Effect of different inputs
Fig. 7 of the main paper shows examples of human-object shapes reconstructed using different combinations of input. Additional results from the configurations analyzed in the second ablation study of Sec. 4.2 and not included in Fig. 7 are shown in Fig. 8. The highest quality human-object shapes are obtained using ReCHOR, confirming what was proved in Sec. 4.2. Smoother shapes are reconstructed when normal maps are excluded (No $S_{N}$) while omitting pose features (No $\sigma_{{h,o}}$) introduces depth ambiguity. Excluding the full-body human input (No $I_{h}$) or the object input (No $I_{o}$) prevents the network from reconstructing the human or object, respectively.
@@ 9.2 Training configurations
ReCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 5: Quantitative results obtained with different training configurations of ReCHOR.
Figure 9: Examples of results obtained with different training configurations of ReCHOR. Front and side views are shown from synHOR in the top row, from BEHAVE [2] in the bottom.
The proposed attention-based neural implicit model is trained end-to-end to enable better joint reasoning about the human-object scene. This section presents results obtained by modifying ReCHOR’s training configuration:
@@
As demonstrated by the quantitative evaluation in Tab. 5 and the qualitative results (Fig. 9), training the attention-based neural implicit model end-to-end significantly improves performance. Both $\mathbf{Sep}$ and $\mathbf{Sep_{no}}$ training configurations result in less accurate human-object reconstructions. Training separate networks for humans and objects leads to reconstruction errors in interaction regions. This demonstrates how the proposed design can embed contextual information about global and local scenes, learning spatial relationship between human and object.
Figure 10: Effect of not applying the inpainting module of ReCHOR before estimating the implicit representation. The same examples used in the paper are illustrated with (w.) and without (w.o) the object.